Use Gauss’s law to find the electric field inside a uniform charged sphere (ρ).


1
Expert's answer
2021-07-05T17:53:43-0400

We know that

Electric flux

"\phi=E.A=EAcos\theta"

"\theta=0°"

"\phi=E.A"

We know that

"E .A=\frac{Q}{\epsilon_0}"

"E(4\pi r^2)=\frac{Q}{\epsilon_0}"

"E_{out}=\frac{Q}{4\pi r^2\epsilon_0}=\frac{kQ}{r^2}" (r>R)

When( r<R)

Total given charge distributed along the sphere surface

Then conducting sphere innerside charge is zero

"E_{in}=0" r<R

When Q=0

"E_{in}=0"